OpenStudy (anonymous):

17. A building lot in a city is shaped as a 30° -60° -90° triangle. The side opposite the 30° angle measures 41 feet. a. Find the length of the side of the lot opposite the 60° angle. b. Find the length of the hypotenuse of the triangular lot. c. Find the sine, cosine, and tangent of the 30° angle in the lot. Write your answers as decimals rounded to four decimal places.

OpenStudy (jdoe0001):



OpenStudy (jdoe0001):



OpenStudy (jdoe0001):

is that correct?

OpenStudy (anonymous):

im not sure that is why i asked. im totally confused

OpenStudy (jdoe0001):

well, that's about the correct diagram for the city lot, now you just need to find the missing legs/sides

OpenStudy (jdoe0001):

to find the bottom leg/side, keep in mind that \(tan(30)=\frac{opp side}{adjacent\ side} \implies adjacent\ side[bottom\ side/leg] = \frac{opp side}{tan(30)}\)

OpenStudy (jdoe0001):

so, \(\frac{41}{tan(30)}\) will give you the bottom side